# Ergotropy transport in a one dimensional spin chain

### v1

@@ I Introduction
The formulation of thermodynamic laws for nanoscale systems has progressed substantially in the last decades Binder et al. (2019); Campbell et al. (2025); Gherardini and De Chiara (2024). From a foundational standpoint, examining how concepts and intuitions must be augmented when the peculiarities of quantum mechanics are considered has provided valuable insights, for example sharpening Landauer’s principle connecting information and energy Reeb and Wolf (2014) and quantum generalisations of the second law Brandão et al. (2015); Strasberg and Winter (2021), among others. Recently, however, there has been a growing need to critically assess the thermodynamics and energetics of emerging quantum technologies in order to ensure their scalability Campbell et al. (2025); Auffèves (2022).
@@
which is the maximal amount of extractable work from the system after undergoing a cyclic unitary process Allahverdyan et al. (2004); Alicki and Fannes (2013). The minimisation is taken over all unitaries, $U_{\tau}$, such that it leaves the state of the system in its lowest available energy configuration, i.e. a passive state $\rho_{p}$. A battery is therefore any system that stores a non-zero amount of extractable work. Viewing a quantum mechanical system as a battery allows us to consider the impact that genuine quantum features, in particular coherence, can have on our understanding of how the laws of thermodynamics must be adapted at the nano-scale Campaioli et al. (2024). Several works have explored the impact of coherence and correlations on ergotropy Binder et al. (2015); Hovhannisyan et al. (2013); Giorgi and Campbell (2015); Francica et al. (2020); Çakmak (2020); Kamin et al. (2020); Fox et al. (2024), designed schemes for stable and efficient charging and discharging protocols Santos et al. (2019); Quach and Munro (2020); Moraes et al. (2022); Barra (2019); Evangelakos et al. (2025); Satriani and Barra (2024); de Moraes et al. (2024); Rinaldi et al. (2025), and explored how different microscopic models behave when modelled as a quantum battery Le et al. (2018); Hoang et al. (2024); Razzoli et al. (2025). Quantum batteries thus provide a useful framework to explore both fundamental aspects of the energetics of quantum systems while also being of potential practical relevance for newly emerging quantum devices Li et al. ; Quach et al. (2022); Tibben et al. (2025). Such advances notwithstanding, recent focus has shifted to exploring the transfer of useful work. This becomes relevant when one considers that the means by which a quantum battery reaches an active state inevitably involves some interaction with a charging system Andolina et al. (2018). It has also been shown that the presence of genuine quantum features can ensure for more efficient transfer of work between quantum battery and a consumption unit Simon et al. (2025).
@@
In this work we complement this line of inquiry by considering a closely related process: the shuttling of extractable work along a spin chain. The use of spin-systems as quantum communication channels is well established Bose (2003, 2007); Kandel et al. (2021); Qiao et al. (2021), with protocols designed to achieve, for example, high-fidelity state transfer and correlation transfer, see e.g. Kay (2010); Christandl et al. (2004, 2005); Bezaz et al. (2025); Alsulami et al. (2024); Nikolopoulos et al. (2014, 2004); Estarellas et al. (2017); Di Franco et al. (2008); Campbell et al. (2011a); Apollaro et al. (2012); Banchi et al. (2011); Vinet and Zhedanov (2012a); Wilkinson et al. (2017). While the information theoretic aspects have been extensively studied, the examination of the thermodynamics of these communication channels is still in its infancy Lewis et al. (2023). The lack of such an analysis may, in part, be due to the fact that for the often considered energy-preserving models that will be the focus of this work, certain properties are communicated identically, e.g. spin or energy. Therefore, one might naively assume that other thermodynamic quantities, such as the ergotropy, would behave similarly. In this work we demonstrate that extractable work is distinct and the ability for a spin chain to effectively transfer ergotropy is delicately dependent on the manner in which it is initially endowed. In particular, we demonstrate that a quantum advantage emerges in the presence of initial coherences. Furthermore, we also assess the thermodynamics associated with switching on the interactions across the chain showing that, while the average work cost can be zero, the nature of the coupling impacts the fluctuations of the work done. In particular, while intrinsically more difficult to engineer, couplings that achieve PST are shown to be more stable at the level of the full work distribution.
@@ II Model
where $\mathcal{J}_{\alpha}\!=\!\frac{1}{2}[(\alpha-1)J+\alpha J_{j,j+1}]$. Due to the excitation preserving nature of the interactions we can restrict to the zero- and single-excitation subspace only. We will therefore find it convenient to work using the basis states $|\underline{\mathbf{1}}\rangle_{n}\!=\!\bigotimes_{j=1}^{n-1}|0\rangle_{j}\otimes|1\rangle_{n}\bigotimes_{l=n+1}^{N}|0\rangle_{l}$ which corresponds to the $n^{\rm th}$ site of the chain in its local excited state, and $|\underline{\mathbf{0}}\rangle\!=\!\bigotimes_{j=1}^{N}|0\rangle_{j}$ corresponding to all spins in the chain to be in their local ground state. The parameter $\alpha$ allows us to interpolate between uniform couplings ($\alpha\!=\!0$) and the couplings required to achieve perfect state transfer (PST) Christandl et al. (2005, 2004) ($\alpha\!=\!1)$. The PST coupling strengths are given by,
@@
For $\alpha\!=\!0$, corresponding to a chain with uniform coupling strengths, the energy eigenvalues take the form Christandl et al. (2005),
@@
In general, uniform couplings do not allow for PST (with the notable exception being extremely small chain sizes of $N\leq 3$). Nevertheless, uniform couplings do allow for some transport. The parameter $J$ sets the overall energy scale for our setup. For $\alpha\in[0,1]$, the width of the energy spectrum of Eq. (2) is bounded in the interval [2J,2J][-2J,2J] centred on, (N2)B-(N-2)B. In Appendices A and B we provide details of the derivations for both limits and refer the interested reader to Refs. Albanese et al. (2004); Christandl et al. (2005); Bernard et al. (2024); Vinet and Zhedanov (2012b) for a more complete analysis.
Ergotropy can be stored in a system in two ways: either through a population inversion and/or in coherences in the energy eigenbasis. When extractable work is present, encoded in either form, the system is said to be in an active state Allahverdyan et al. (2004). Our interest will be on demonstrating that initially storing ergotropy in quantum coherences is highly advantageous with respect to the ability to shuttle it along the chain when the couplings are not identically those which achieve PST. We therefore consider the following initial states for the chain,
@@
In Eq. (9) we see the first site is initially pure and, therefore, the initial ergotropy to be transported is endowed in the coherence. In contrast, the first site in Eq. (10) is a mixed state in its local energy eigenbasis. Ergotropy can be encoded in this state also, however it requires a population inversion, i.e. q>0.5q\!>\!0.5. To ensure a fair comparison, we will consider states which contain the same amount of initial ergotropy which, by directly computing Eq. (1) for the first site of the chain, corresponds to the condition $q\!=\!\left[1+\sin^{2}\left(\frac{\theta}{2}\right)\right]/2$. Finally, we remark that although $\phi$ in Eq. (9), plays a significant role in a system’s dynamics, e.g fidelity, it does not play a role on the level of the energetics Bayat et al. (2011). Therefore, without loss of generality we set $\phi=0$.
@@ III.1 Evolved state and transition probabilities
The choice of initial states in Eqs. (9) and (10) and excitation preserving form of Eq. (2) means that we can restrict our analysis to a zero- and single-excitation subspace of the total Hamiltonian. The time dependent reduced density matrix for the $n^{\rm th}$ spin in the chain is analytically attainable and takes the convenient form Campbell et al. (2011b),
@@
As our focus is on examining the locally extractable work that is transported along the chain, we assume the following protocol. At t=0t\!=\!0 the interactions are suddenly quenched on and the whole system evolves for a time, $t$. The system evolves up to the first reflection time, at which point we apply a swap operation between the last site of the chain and an auxiliary qubit with local Hamiltonian $H\!=\!-B\sigma_{z}$. We compute the ergotropy of this auxiliary system using Eq. (1) which exactly corresponds to the ergotropy tranferred via the spin chain. It should be immediately clear that for $\alpha\!=\!1$ the PST couplings guarantee that whatever state is initially at site 1 will perfectly reach the end of the chain given a suitable amount of time, $Jt\!=\!\frac{\pi}{4\,G_{N}}N$ Christandl et al. (2004), (and in fact this state will continue to bounce back and forth along the chain precisely localising at either end in a perfectly periodic manner). Thus, the strict limit of $\alpha\!=\!1$ guarantees the perfect shuttling of ergotropy, or in fact any quantity encoded in the initial state of the first site. However, this necessitates that the precise distribution of PST couplings is realised. We will therefore focus on $\alpha\!<\!1$, and in particular the uniform coupling limit $\alpha\!=\!0$, which will provide a lower bound for the ergotropy transport facilitated by our model. Furthermore, this focus will allow us to demonstrate that, in general, coherently endowed ergotropy can be more efficiently transported.
@@ III.2 Ergotropy transport for non-PST coupling spin chains
where the constraint follows since ergotropy is only non-zero for an active state. We begin by examining how effectively the chain is able to transport the extractable work for various chain lengths and amounts of initially stored ergotropy. For simplicity we focus on the first reflection time which (generally) corresponds to the maximum transport efficiency,
@@
and determine the value of ergotropy that can be extracted from the end of the chain at site $N$ at this time, which we denote as $\varepsilon_{max}$. Eq. (16) in the PST limit is known Christandl et al. (2004), and we numerically determine the quadratic form in $\alpha$ that interpolates between the two limits.
@@
Fig. 1(a) demonstrates that for a mixed state with a given $\varepsilon_{in}$ there exists a strict chain-length cut-off beyond which no more work can be locally extracted. In contrast, for coherence endowed ergotropy, $\varepsilon_{max}$ scales to leading order as $\sim N^{-2/3}$. This scaling stems from the transition amplitude function in Eq. (13), which for a uniformly coupled chain scales as $f_{N}(T)\sim N^{-1/3}$ Bose (2007) for large $N$ and sufficiently long times. This scaling can be seen by examining the thermodynamic limit of Eq. (13), which we derive in Appendix B and is discussed further in Ref. Bose (2007). From Eq. (14), we see that $\varepsilon_{coh}(T)$, to leading order, scales as $|f_{N}(T)|^{2}\!\propto\!N^{-2/3}$. Thus, for a finite system with coherences, ergotropy will reach the end of the chain given sufficient time. In panels (b) and (c) we show that the efficiency of ergotropy transport increases as we tune $\alpha$. For mixed initial states the maximum length of chain length for which ergotropy can still be transported increases and, as was the case for uniform couplings, the larger the amount of initial ergotropy present, the further it can be transported. For ergotropy initially endowed in coherences, increasing $\alpha$ has a similar impact insofar as the efficiency of the transport improves with increasing $\alpha$. For the moderate size systems shown here we see that the more initial ergotropy inputted corresponds to a larger amount reaching the end of the chain. While this is true in absolute terms for chains up to $N\sim 100$ as shown, we remark that the dashed curves can cross for sufficiently large chains. Furthermore, if we consider the relative ergotropy transported, i.e. the efficiency $\varepsilon_{max}/\varepsilon_{in}$, a lower amount of initial ergotropy tends to be more faithfully transferred. This demonstrates that it is more efficient to transfer extractable work across long chains via coherences; the only exception to this being the strict limit of $\alpha\!=\!1$ where PST is achieved and therefore the ergotropy will be faithfully transferred regardless of how it is initially encoded.
@@
Figure 2: $\varepsilon_{max}$ endowed in coherences (dashed curves) and populations (solid curves) of the last qubit in the chain evaluated at JtJt given by Eq. (16), as a function of $\varepsilon_{in}$. We show three system sizes, $N=\{10,25,100\}$. Parameters: $\alpha=0,J=1,B=1$. Inset: Peak value of $\varepsilon_{max}$ as a function of chain size NN. The corresponding value is marked with a star for the system sizes shown in the main panel.
@@
We further consider the efficiency of the ergotropy transport in Fig. 2 where we show the extractable work at the last site of the chain, $\varepsilon_{max}$, as a function of the initial ergotropy, $\varepsilon_{in}$, for the uniform coupling case. The dashed curves show the behavior for Eq. (9), i.e. an active pure state. For small chains, $N\sim 10$, increasing $\varepsilon_{in}$ leads to a greater amount of ergotropy transported. However, for larger chains we see this no longer holds. As $N$ is increased, we find the optimal initial state to ensure maximum transport of useful work tends to an input state that is an equal coherent superposition on the first site. In contrast, for a mixed state, the higher the population of the excited state, the more ergotropy will reach the end of the chain, however, again we find that there is a strict cut off in chain length above which no useful work can be transferred. In the inset of Fig. 2 we plot the peak of each of the dashed curves as a function of chain length, $N$, from which we see that in the thermodynamic limit it is most favourable to initialise the state of the first qubit in an equal superposition. We show this analytically using the thermodynamic limit expression of Eq. (13) in Appendix B.

### v2

@@ I Introduction
The formulation of thermodynamic laws for nanoscale systems has progressed substantially in the last decades Binder2019book; Campbell2025roadmap; Gherardini2024. From a foundational standpoint, examining how concepts and intuitions must be augmented when the peculiarities of quantum mechanics are considered has provided valuable insights, for example sharpening Landauer’s principle connecting information and energy Reeb2014 and quantum generalisations of the second law Brand_o2015; Strasberg2021, among others. Recently, however, there has been a growing need to critically assess the thermodynamics and energetics of emerging quantum technologies in order to ensure their scalability Campbell2025roadmap; Auffeves2022.
@@
which is the maximal amount of extractable work from the system after undergoing a cyclic unitary process Allahverdyan2004; Alicki2013PRE. The minimisation is taken over all unitaries, $U_{\tau}$, such that it leaves the state of the system in its lowest available energy configuration, i.e. a passive state $\rho_{p}$. A battery is therefore any system that stores a non-zero amount of extractable work. Viewing a quantum mechanical system as a battery allows us to consider the impact that genuine quantum features, in particular coherence, can have on our understanding of how the laws of thermodynamics must be adapted at the nano-scale Campaioli2024Colloquium. Several works have explored the impact of coherence and correlations on ergotropy binder2015quantacell; Hovhannisyan2013PRL; giorgi2015correlation; QCohErg; Cakmak2020PRE; Kamin2020PRE; NthRoot; mondal2025; Tirone2021, designed schemes for stable and efficient charging and discharging protocols Santos2019PRE; Quach2020PRApp; moraes2022charging; Barra20219PRL; evangelakos2025rapid; satriani2024daemonic; deMoraes2024quantum; Rinaldi2025PRA; Tirone2023, and explored how different microscopic models behave when modelled as a quantum battery Le2018Spin; Hoang2024; razzoli2025cyclic. Quantum batteries thus provide a useful framework to explore both fundamental aspects of the energetics of quantum systems while also being of potential practical relevance for newly emerging quantum devices Li2025; quach2022superabsorption; tibben2025PRXEnergy. Such advances notwithstanding, recent focus has shifted to exploring the transfer of useful work. This becomes relevant when one considers that the means by which a quantum battery reaches an active state inevitably involves some interaction with a charging system Andolina2018PRB. It has also been shown that the presence of genuine quantum features can ensure for more efficient transfer of work between quantum battery and a consumption unit CorrLossErg.
@@
In this work we complement this line of inquiry by considering a closely related process: the shuttling of extractable work along a spin chain. The use of spin-systems as quantum communication channels is well established QcommUnmod; QcommReview; Kandel2021; Qiao2021, with protocols designed to achieve, for example, high-fidelity state transfer and correlation transfer, see e.g. kay2010perfect; PSTQCA; PSTarbQCA; bezaz2025quasi; alsulami2024scalable; nikolopoulos2014quantum; nikolopoulos2004electron; Estarellas2017PRA; DiFranco2008PRL; Campbell2011PRA; Apollaro2012PRA; banchi2011long; Vinet2012PRA; wilkinson2017rapid. While the information theoretic aspects have been extensively studied, the examination of the thermodynamics of these communication channels is still in its infancy LewisPRB2023. The lack of such an analysis may, in part, be due to the fact that for the often considered energy-preserving models that will be the focus of this work, certain properties are communicated identically, e.g. spin or energy. Therefore, one might naively assume that other thermodynamic quantities, such as the ergotropy, would behave similarly. In this work we demonstrate that extractable work is distinct and the ability for a spin chain to effectively transfer ergotropy is delicately dependent on the manner in which it is initially endowed. In particular, we demonstrate that a quantum advantage emerges in the presence of initial coherences. Furthermore, we also assess the thermodynamics associated with switching on the interactions across the chain showing that, while the average work cost can be zero, the nature of the coupling impacts the fluctuations of the work done. In particular, while intrinsically more difficult to engineer, couplings that achieve PST are shown to be more stable at the level of the full work distribution.
@@ II Model
where $\mathcal{J}_{\alpha}\!=\!\frac{1}{2}[(\alpha-1)J+\alpha J_{j,j+1}]$. Due to the excitation preserving nature of the interactions we can restrict to the zero- and single-excitation subspace only. We will therefore find it convenient to work using the basis states $|\underline{\mathbf{1}}\rangle_{n}\!=\!\bigotimes_{j=1}^{n-1}|0\rangle_{j}\otimes|1\rangle_{n}\bigotimes_{l=n+1}^{N}|0\rangle_{l}$ which corresponds to the $n^{\rm th}$ site of the chain in its local excited state, and $|\underline{\mathbf{0}}\rangle\!=\!\bigotimes_{j=1}^{N}|0\rangle_{j}$ corresponding to all spins in the chain to be in their local ground state. The parameter $\alpha$ allows us to interpolate between uniform couplings ($\alpha\!=\!0$) and the couplings required to achieve perfect state transfer (PST) PSTarbQCA; PSTQCA ($\alpha\!=\!1)$. The PST coupling strengths are given by,
@@
For $\alpha\!=\!0$, corresponding to a chain with uniform coupling strengths, the energy eigenvalues take the form PSTarbQCA,
@@
In general, uniform couplings do not allow for PST (with the notable exception being extremely small chain sizes of $N\leq 3$). Nevertheless, uniform couplings do allow for some transport. The parameter $J$ sets the overall energy scale for our setup. For $\alpha\in[0,1]$, the width of the energy spectrum of Eq. (2) is bounded in the interval [2J,2J][-2J,2J] centred on, (N2)B-(N-2)B. In Appendices A and B we provide details of the derivations for both limits and refer the interested reader to Refs. Albanese2004; PSTarbQCA; inhomspinchain; Vinet2012PRA for a more complete analysis.
Ergotropy can be stored in a system in two ways: either through a population inversion and/or in coherences in the energy eigenbasis. When extractable work is present, encoded in either form, the system is said to be in an active state Allahverdyan2004. Our interest will be on demonstrating that initially storing ergotropy in quantum coherences is highly advantageous with respect to the ability to shuttle it along the chain when the couplings are not identically those which achieve PST. We therefore consider the following initial states for the chain,
@@
In Eq. (9) we see the first site is initially pure and, therefore, the initial ergotropy to be transported is endowed in both the populations and coherence. In contrast, the first site in Eq. (10) is a mixed state in its local energy eigenbasis. Ergotropy can be encoded in this state also, however it requires a population inversion, i.e. q>0.5q\!>\!0.5. To ensure a fair comparison, we will consider states which contain the same amount of initial ergotropy which, by directly computing Eq. (1) for the first site of the chain, corresponds to the condition $q\!=\!\left[1+\sin^{2}\left(\frac{\theta}{2}\right)\right]/2$. Finally, we remark that although $\phi$ in Eq. (9), plays a significant role in a system’s dynamics, e.g fidelity, it does not play a role on the level of the energetics Bayat2011. Therefore, without loss of generality we set $\phi=0$.
@@ III.1 Evolved state and transition probabilities
The choice of initial states in Eqs. (9) and (10) and excitation preserving form of Eq. (2) means that we can restrict our analysis to a zero- and single-excitation subspace of the total Hamiltonian. The time dependent reduced density matrix for the $n^{\rm th}$ spin in the chain is analytically attainable and takes the convenient form Campbell2011PRA,
@@
As our focus is on examining the locally extractable work that is transported along the chain, we assume the following protocol. At t=0t\!=\!0 the interactions are suddenly quenched on and the whole system evolves for a time, $t$. The system evolves up to the first reflection time, i.e. the first time at which the transfer function $f_{N}(t)$ reaches a local maximum, at which point we apply a swap operation between the last site of the chain and an auxiliary qubit with local Hamiltonian $H\!=\!-B\sigma_{z}$. We compute the ergotropy of this auxiliary system using Eq. (1) which exactly corresponds to the ergotropy tranferred via the spin chain. It should be immediately clear that for $\alpha\!=\!1$ the PST couplings guarantee that whatever state is initially at site 1 will perfectly reach the end of the chain given a suitable amount of time, $Jt\!=\!\frac{\pi}{4\,G_{N}}N$ PSTQCA, (and in fact this state will continue to bounce back and forth along the chain precisely localising at either end in a perfectly periodic manner). Thus, the strict limit of $\alpha\!=\!1$ guarantees the perfect shuttling of ergotropy, or in fact any quantity encoded in the initial state of the first site. However, this necessitates that the precise distribution of PST couplings is realised. We will therefore focus on $\alpha\!<\!1$, and in particular the uniform coupling limit $\alpha\!=\!0$, which will provide a lower bound for the ergotropy transport facilitated by our model. Furthermore, this focus will allow us to demonstrate that, in general, coherently endowed ergotropy can be more efficiently transported.
@@ III.2 Ergotropy transport for non-PST coupling spin chains
where the constraint follows since ergotropy is only non-zero for an active state. We begin by examining how effectively the chain is able to transport the extractable work for various chain lengths and amounts of initially stored ergotropy. For simplicity we focus on the first reflection time which (generally) corresponds to the maximum ergotropy transport,
@@
and determine the value of ergotropy that can be extracted from the end of the chain at site $N$ at this time, which we denote as $\varepsilon_{max}$. Eq. (16) in the PST limit is known PSTQCA, and we numerically determine the quadratic form in $\alpha$ that interpolates between the two limits. We note that from the preceding, it follows that other sites in the chain may host extractable work during the dynamics. For initial states of the form Eq. (10) we find that the conditions necessary for most of the sites in the bulk of the chain to become active are not met, i.e. they do not achieve a population inversion. In contrast, initialising the chain in Eq. (9) we find that the presence of coherence at a given site leads to the possibility of ergotropy being dynamically present. Nevertheless, in both cases, it is the hard-boundary that causes a “refocussing” of the ergotropy at the last site of the chain and will be our primary focus.
@@
Fig. 1(a) demonstrates that for a mixed state with a given $\varepsilon_{in}$ there exists a strict chain-length cut-off beyond which no more work can be locally extracted. In contrast, for coherence endowed ergotropy, $\varepsilon_{max}$ scales to leading order as $\sim N^{-2/3}$. This scaling stems from the transition amplitude function in Eq. (13), which for a uniformly coupled chain scales as $f_{N}(T)\sim N^{-1/3}$ QcommReview for large $N$ and sufficiently long times. This scaling can be seen by examining the thermodynamic limit of Eq. (13), which we derive in Appendix B and is discussed further in Ref. QcommReview. From Eq. (14), we see that $\varepsilon_{coh}(T)$, to leading order, scales as $|f_{N}(T)|^{2}\!\propto\!N^{-2/3}$. Thus, for a finite system with coherences, ergotropy will reach the end of the chain given sufficient time. In panels (b) and (c) we show that more ergotropy can be faithfully transported as we tune $\alpha$. For mixed initial states the maximum length of chain length for which ergotropy can still be transported increases and, as was the case for uniform couplings, the larger the amount of initial ergotropy present, the further it can be transported. For ergotropy initially endowed in coherences, increasing $\alpha$ has a similar impact insofar as the efficiency of the transport improves with increasing $\alpha$. For the moderate size systems shown here we see that the more initial ergotropy inputted corresponds to a larger amount reaching the end of the chain. While this is true in absolute terms for chains up to $N\sim 100$ as shown, we remark that the dashed curves can cross for sufficiently large chains. Furthermore, if we consider the relative ergotropy transported, i.e. the efficiency $\varepsilon_{max}/\varepsilon_{in}$, a lower amount of initial ergotropy tends to be more faithfully transferred. This demonstrates that it is more efficient to transfer extractable work across long chains via coherences; the only exception to this being the strict limit of $\alpha\!=\!1$ where PST is achieved and therefore the ergotropy will be faithfully transferred regardless of how it is initially encoded.
@@
Figure 2: $\varepsilon_{max}$ endowed in coherences (dashed curves) and populations (solid curves) of the last qubit in the chain evaluated at JtJt given by Eq. (16), as a function of $\varepsilon_{in}$. We show three system sizes, $N=\{10,25,100\}$. Parameters: $\alpha=0,J=1,B=1$. Inset: The value of $\varepsilon_{in}$ that corresponds to the peak value of $\varepsilon_{max}$, as a function of chain size NN. The corresponding value is marked with a star for the system sizes shown in the main panel.
@@
We further consider the efficiency of the ergotropy transport in Fig. 2 where we show the extractable work at the last site of the chain, $\varepsilon_{max}$, as a function of the initial ergotropy, $\varepsilon_{in}$, for the uniform coupling case. The dashed curves show the behavior for Eq. (9), i.e. an active pure state. For small chains, $N\sim 10$, increasing $\varepsilon_{in}$ leads to a greater amount of ergotropy transported. However, for larger chains we see this no longer holds. As $N$ is increased, we find the optimal initial state to ensure maximum transport of useful work tends to an input state that is an equal coherent superposition on the first site. Thus, for initial states of the form Eq. (9) and finite chains, there is a non-trivial interplay between the role of coherence in facilitating ergotropy transport over longer chains and the benefits of having more initial ergotropy which necessitates a population inversion. In contrast, for a mixed state, the higher the population of the excited state, the more ergotropy will reach the end of the chain, however, again we find that there is a strict cut off in chain length above which no useful work can be transferred. In the inset of Fig. 2 we plot the the value of $\varepsilon_{in}$, labelled $\varepsilon_{\star}$, that corresponds to the peak of each of the dashed curves as a function of chain length, $N$, from which we see that in the thermodynamic limit it is most favourable to initialise the state of the first qubit in an equal superposition. We show this analytically using the thermodynamic limit expression of Eq. (13) in Appendix B.